OpenStudy (anonymous):

This problem only has 1 value, so I don't really understand how it is possible to solve. Thanks for any help! A particular color of light has a wavelength of 550 nm traveling in water. Calculate the wavelength in air. I know the answer is 732 nm, but I don't get why. Thanks for the help!

OpenStudy (anonymous):

@hartnn @dan815 @zepdrix @jdoe0001 @Loser66 @Luigi0210 @mathmale @Mehek14 @jhonyy9 @jigglypuff314 @agent0smith @jabez177

OpenStudy (agent0smith):

The frequency of light is the same in different media. The speed and wavelength vary. It seems like you need to know (or are meant to use given values) for the speed in water and air. In air you can use 3x10^8 but water is slower.

OpenStudy (agent0smith):

\[\large v = f \lambda \] solve for f, since it's equal in water and air, \[\large f= \frac{ v }{ \lambda }\] then you can use \[\large \frac{ v_1 }{ \lambda_1 }=\frac{ v_2 }{\lambda_2 }\]

OpenStudy (anonymous):

oh I see! Thanks!

OpenStudy (anonymous):

Yup, I just solved it out. I got 733.33, but I'm guessing my teacher rounded for his answer. Thanks a lot!